Geometric Distribution (Optional) Quiz
The question sheet
Reveal any answer as you study-
How many main characteristics does a geometric experiment have?
- Three
- Two
- Five
- Four
Reveal answer
Answer: Three
Source evidence
PDF page 262: There are three main characteristics of a geometric experiment:
-
In a geometric experiment, you keep repeating a trial until:
- Every trial succeeds
- Two failures occur
- A fixed number of trials
- The first success, then stop
Reveal answer
Answer: The first success, then stop
Source evidence
PDF page 262: experiment with number of trials n = 1. In other words, you keep repeating what you are doing until the first success. Then you stop. For example, you throw a dart at a bull's-eye until you hit the bull's-eye. The first time you hit the bull's-eye is a success so you stop throwing the dart. It might take six tries until you hit the bull's-eye. You can think of the trials as failure, failure, failure, failure, failure, success, stop.
-
For a geometric random variable, X is defined as:
- Number of failures total
- Number of successes in n trials
- Probability of a success
- Number of trials until first success
Reveal answer
Answer: Number of trials until first success
Source evidence
PDF page 263: X = the number of independent trials until the first success.
-
In a geometric distribution, q represents:
- The mean of X
- 1 - p, the probability of a failure
- The probability of a success
- The number of trials
Reveal answer
Answer: 1 - p, the probability of a failure
Source evidence
PDF page 263: p = the probability of a success, q = 1 – p = the probability of a failure.
-
In the win/lose game where p = 0.57, the probability of a failure q equals:
- 1.00
- 0.57
- 0.43
- 0.17
Reveal answer
Answer: 0.43
Source evidence
PDF page 263: Suppose a game has two outcomes, win or lose. You repeatedly play that game until you lose. The probability of losing is p = 0.57. If we let X = the number of games you play until you lose (includes the losing game), then X is a geometric random variable. All three characteristics are met. Each game you play is a Bernoulli trial, either win or lose. You would need to play at least one game before you stop. X takes on the values 1, 2, 3, . . . (could go on indefinitely). Since we are measuring the number of games you play until you lose, we define a success as losing a game and a failure as winning a game. The probability of a success
PDF page 263: and the probability of a failure q = 1 – p = 1 – 0.57 = 0.43. Both p and q remain
-
In the win/lose game example, a success is defined as:
- Winning a game
- Losing a game
- Stopping the game
- Playing a game
Reveal answer
Answer: Losing a game
Source evidence
PDF page 263: Suppose a game has two outcomes, win or lose. You repeatedly play that game until you lose. The probability of losing is p = 0.57. If we let X = the number of games you play until you lose (includes the losing game), then X is a geometric random variable. All three characteristics are met. Each game you play is a Bernoulli trial, either win or lose. You would need to play at least one game before you stop. X takes on the values 1, 2, 3, . . . (could go on indefinitely). Since we are measuring the number of games you play until you lose, we define a success as losing a game and a failure as winning a game. The probability of a success
-
In the win/lose game, X takes on the values:
- 0, 1, 2, 3
- Only 1
- 0 to 100
- 1, 2, 3, . . .
Reveal answer
Answer: 1, 2, 3, . . .
Source evidence
PDF page 263: Suppose a game has two outcomes, win or lose. You repeatedly play that game until you lose. The probability of losing is p = 0.57. If we let X = the number of games you play until you lose (includes the losing game), then X is a geometric random variable. All three characteristics are met. Each game you play is a Bernoulli trial, either win or lose. You would need to play at least one game before you stop. X takes on the values 1, 2, 3, . . . (could go on indefinitely). Since we are measuring the number of games you play until you lose, we define a success as losing a game and a failure as winning a game. The probability of a success
-
In the safety engineer example, the probability of a success p equals:
- 0.02
- 0.35
- 0.65
- 0.17
Reveal answer
Answer: 0.35
Source evidence
PDF page 263: A safety engineer feels that 35 percent of all industrial accidents in her plant are caused by failure of employees to follow instructions. She decides to look at the accident reports (selected randomly and replaced in the pile after reading) until she finds one that shows an accident caused by failure of employees to follow instructions. If we let X = the number of accidents the safety engineer must examine until she finds a report showing an accident caused by employee failure to follow instructions, then X is a geometric random variable. All three characteristics are met. Each accident report she reads is a Bernoulli trial: the accident was either caused by failure of employees to follow instructions or not. She would need to read at least one accident report before she stops. X takes on the values 1, 2, 3, . . . (could go on indefinitely). Since we are measuring the number of reports she needs to read until one that shows an accident caused by failure of employees to follow instructions, we define a success as an accident caused by failure of employees to follow instructions. If an accident was caused by another reason, the report is defined as a failure. The probability of a success p = .35 and the probability of a failure . Both p and q remain the same from report to report. If we want to find the probability that the safety engineer will have to examine at least three reports until she finds a report showing an accident caused by employee failure to follow instructions, then the probability could be written as . If we want to find how many reports, on average, the safety engineer would expect to look at until she finds a report showing an accident caused by employee failure to follow instructions, we need to find the expected value E(x). We will explain how to solve these questions later in this section.
-
In the dartboard example, your probability of hitting the center p equals:
- 0.10
- 0.57
- 0.35
- 0.17
Reveal answer
Answer: 0.17
Source evidence
PDF page 263: You throw darts at a board until you hit the center area. Your probability of hitting the center area is p = 0.17. You want to find the probability that it takes eight throws until you hit the center. What values does X take on?
-
In the special printer ink example, the probability of success p equals:
- 35 percent
- 10 percent
- 55 percent
- 15 percent
Reveal answer
Answer: 10 percent
Source evidence
PDF page 264: You need to find a store that carries a special printer ink. You know that of the stores that carry printer ink, 10 percent of them carry the special ink. You randomly call each store until one has the ink you need. What are p and q?
-
The notation X ~ G(p) is read as:
- X has a binomial distribution
- X has a uniform distribution
- X has a geometric distribution
- X has a normal distribution
Reveal answer
Answer: X has a geometric distribution
Source evidence
PDF page 264: X ~ G(p) Read this as X is a random variable with a geometric distribution. The parameter is p; p = the probability of a success for each trial.
-
For the defective component example, the value of p is:
- 0.55
- 0.35
- 0.07
- 0.02
Reveal answer
Answer: 0.02
Source evidence
PDF page 265: probability that the first defect is caused by the seventh component tested. How many components do you expect to test until one is found to be defective? Let X = the number of computer components tested until the first defect is found. X takes on the values 1, 2, 3, . . . where p = .02. X ~ G(.02) Find P(x = 7). There is a formula to define the probability of a geometric distribution . We can use the formula to find . But since the calculation is tedious and time consuming, people usually use a graphing calculator or software to get the answer. Using a graphing calculator, you can get . The instruction of TI83, 83+, 84, 84+ is given below.
Statistics
Statistics by OpenStax, used under CC BY 4.0. Changes made by Stratacademy.
Make your own — free
Turn any notes into a game in under a minute. Free to start.